) Find the area under the standard normal curve between 0.15 and 1.6.
answer:

(b) Find the area under the standard normal curve between -1.55 and 1.6.
answer:

(b) Find the area under the standard normal curve between -1.55 and -0.15.
answer:
Answer #1

Solution :

(a)

(a)
P(0.15 < z < 1.6) = P(z < 1.6) - P(z < 0.15) = 0.3856

Area = 0.3856

(b)

P(-1.55 < z < 1.6) = P(z < 1.6) - P(z < -1.55) = 0.8846

Area = 0.8846

(c)

P(-1.55 < z < -0.15) = P(z < -0.15) - P(z < -1.55) = 0.3798

Area = 0.3798